Let R be a relation on ℤ given by xRy if and only if x²-y² is divisible by 3. Show that this relation is an equivalence relation and find its corresponding equivalence classes.
1
Expert's answer
2020-02-21T10:18:10-0500

First we show that given relation R is an equivalence relation:

"1.\ \forall x\ xRx: \quad 3 \mid 0 \implies 3 \mid x^2 - x^2"

"2.\ xRy \implies yRx : \quad 3 \mid x^2 - y^2 \implies 3 \mid-(x^2 - y^2) \implies 3 \mid y^2 - x^2"

"3.\ xRy, yRz \implies xRz: \ 3 \mid x^2 - y^2 ,\ 3 \mid y^2 - z^2 \implies \\ \implies 3 \mid (x^2 - y^2) + (y^2 - z^2) \implies 3 \mid x^2 - z^2"

Hence by definition it is indeed an equivalence relation.

Now we find the equivalence classes:

"3 \mid x^2 - y^2 \\
3 \mid (x-y)(x+y) \\
3 \mid x - y \quad or \quad 3 \mid x+y"

"x = y + 3k \quad or \quad x = -y + 3k, \quad k \in \mathbb{Z}"

We can see from here, that for every "y \in \mathbb{Z}" the corresponding equivalence class is:

"[y] = \{ x \in \mathbb{Z} \mid x = \pm y + 3k, \ k \in \mathbb{Z} \}"